# Seeing the Forest for the Trees: A Large Scale, Continuously Updating Meta-Analysis of Frontier LLMs

### v1

@@ Abstract
The surge of LLM studies makes synthesizing their findings challenging. Meta-analysis can uncover important trends across studies, but its use is limited by the time-consuming nature of manual data extraction. Our study presents a semi-automated approach for meta-analysis that accelerates data extraction using LLMs. It automatically identifies relevant arXiv papers, extracts experimental results and related attributes, and organizes them into a structured dataset. We conduct a comprehensive meta-analysis of frontier LLMs using an automatically extracted dataset, reducing the effort of paper surveying and data extraction by more than 93% compared to manual approaches. We validate our dataset by showing that it reproduces key findings from a recent manual meta-analysis about Chain-of-Thought (CoT), and also uncovers new insights that go beyond it, showing for example that in-context examples benefit multimodal tasks but offer limited gains in mathematical tasks compared to CoT. Our automatically updatable dataset enables continuous tracking of target models by extracting evaluation studies as new data becomes available. Through our scientific artifacts and empirical analysis, we provide novel insights into LLMs while facilitating ongoing meta-analyses of their behavior. [^1] Our data and code are available in https://github.com/JJumSSu/meta-analysis-frontier-LLMs
Seeing the Forest for the Trees:
A Large Scale, Continuously Updating Meta-Analysis of Frontier LLMs
@@ 1 Introduction
Figure 1: The diagram of our semi-automated meta-analysis process and a key finding derived from data automatically extracted from the arXiv database.
@@
The rapid advancement of Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; Anthropic, 2024) has led to a proliferation of empirical studies. Recent “surveys of surveys” (Lv et al., 2024; Ruan et al., 2024) indicates the overwhelming growth of LLM research publications, surpassing what individual researchers can manually review. The growing number of evaluations—each using different models, datasets, and prompting configurations—makes it challenging to analyze and synthesize findings across studies, underscoring the urgent need for systematic analysis.
@@
Meta-analysis, which examines data from multiple independent studies to identify overall trends, offers a solution (Reiter, 2018; Hupkes et al., 2023). However, conducting such analysis is time-consuming (Yun et al., 2023, 2024). For instance, as shown in Fig. 1, understanding the effects of combining Chain-of-Thought (CoT) with in-context examples requires identifying relevant papers, extracting specific experimental results, and aggregating the data. Furthermore, meta-analyses are likely to become quickly outdated, in a fast-moving field, motivating the need for automated analyses that can be continuously updated as new results are shared on preprint servers such as arXiv.org.
@@
In this study, we conduct the most extensive quantitative meta-analysis of frontier LLMs to date. Our dataset comprises 18,127 experimental records extracted from 1,737 papers, along with 359 referenced papers describing datasets used in these experiments. Our dataset can also be dynamically updated with minimal effort to incorporate results from new models or studies as they are published. This comprehensive dataset enables us to present new insights into LLM performance, a first in the field where previous studies on extracting leaderboards from machine learning papers have primarily focused on improving data extraction accuracy (Kardas et al., 2020; Yun et al., 2024).
@@
To achieve this, we experiment with an automated approach to data extraction that efficiently processes research literature. This approach uses an LLM to scan the arXiv database, identify relevant papers with experimental results on target models, and extract experimental results along with pertinent attributes. This reduces manual effort needed for surveying and extraction by more than 93% compared to expert manual data extraction. The approach employs schema-driven extraction (Bai et al., 2023) and context augmentation from paper contents to capture essential attributes for a comprehensive understanding of evaluation results. Additionally, the process creates dataset descriptions that summarize key characteristics of the dataset associated with the evaluation record, enriching the depth of dataset utility in meta-analysis.
@@
Our analysis shows that our dataset can efficiently replicate a previous manual meta-analysis (Sprague et al., 2024), which found that CoT reasoning primarily benefits mathematical and symbolic reasoning tasks. We also identify three novel insights from the dataset on prompting configurations. First, in-context examples enhance multimodal tasks but offer limited benefits for mathematical tasks compared to a zero-shot setup. Second, our qualitative analysis identifies certain dataset characteristics linked to negative impacts when using CoT or in-context learning (ICL). Lastly, CoT with demonstrations (i.e., ICL) generally outperforms CoT without them. However, CoT’s improvement over direct (i.e. standard) prompting stays consistent with the same number of demonstrations, whether zero-shot or few-shot. To further contribute to the understanding of frontier LLMs’ capabilities, we will open-source our dataset and code, streamlining meta-analyses with automated and continuous study integration.
@@ 2 Data Extraction
Automated data extraction is essential for scaling and improving meta-analysis efficiency. This section details our pipeline for extracting experimental results and model attributes from arXiv sources. We define target models and attributes (section 2.1) and introduce our LLM-powered extraction process, which includes three stages: preprocessing & filtering (section 2.2), extraction & augmentation (section 2.3), and description generation (section 2.4). We utilize GPT-4o (Hurst et al., 2024) throughout the pipeline.
@@ 2.1 Defining Target Models and Attributes
To enable thorough analysis across different studies, we extract various performance-related fields from proprietary models. The targeted attributes include: Dataset Name, Subset, Model Name, Prompting Method, Number of In-Context Examples (shots), Metric Name, and Performance. Note that the subset refers to any division of the dataset, such as a subtask, domain, split, or language pair. For instance, as shown in Fig.2, the arXiv paper with ID 2301.08721 (Cheng et al., 2023) presents experimental results for GPT-4 (Achiam et al., 2023) on the SVAMP dataset (Patel et al., 2021). The evaluation used Batch Prompting with CoT (Wei et al., 2022; Cheng et al., 2023) and 12 in-context samples (Brown et al., 2020), achieving an accuracy score of 95.0.
@@ 2.2 Preprocessing & Filtering
To extract experimental results of target models and attributes, we used arXiv sources from January 2023 to December 2024, focusing on machine learning papers (cs.CV, cs.AI, cs.CL, cs.LG). We downloaded LaTeX source files and applied regex-based methods to extract structured data like tables (Bai et al., 2023). Before extraction, we filtered tables to reduce time and API costs by selecting those with target model results using keywords like “gemini” and “gpt”. We concentrated on leaderboard tables presenting main benchmark results, as described in Kardas et al. (2020), and used Llama3.1-70B (Dubey et al., 2024) for cost-efficient filtering.
@@ 2.3 Extraction & Augmentation
2.4 Dataset Description Generation
In addition to extracting structured representations of experiments, we also generate summary descriptions of the relevant tasks and datasets (Fig. 3) that can aid in-depth meta-analysis, such as classifying records by research subarea. We use a two-stage approach to create these descriptions. Initially, the LLM generates descriptions using its internal knowledge based on the dataset name and subset, which is cost-effective. When the LLM is uncertain, especially with lesser-known datasets or due to a knowledge cut-off, we prompt it to extract descriptions using the full content of the source papers as a reference. The source papers can be either the paper containing the table or the original dataset paper cited within it, which is retrieved using the BibTex references obtained from section 2.3.
@@ 3 Comprehensive Analysis of the Dataset
Table 1 presents sampled instances of our dataset, while the dataset’s statistics are summarized in Table 2. A total of 18,127 experimental records of four target models were extracted from scanning over 300k arXiv source papers. GPT-4 (Achiam et al., 2023) and GPT-4o (Achiam et al., 2023) dominate the results, while Claude 3 Opus (Anthropic, 2024) and Gemini 1.0 Pro (Team et al., 2023) have significantly fewer entries, highlighting a bias favoring certain proprietary models. Many prompting configurations have missing values, confirmed by human verification (section 3.1) as unreported by the original authors, likely indicating the default setup (standard prompting & zero-shot). The majority of dataset descriptions are generated referencing the source paper of the table, as these papers often describe datasets when presenting benchmarks or detailing experimental sections.
@@
3.1 Quality and Efficiency Assessment
Our human evaluation assessed extraction accuracy and description quality across 40 records (280 total attributes) from different papers. Two NLP Ph.D. students verified field extraction correctness (Bai et al., 2023) and rated descriptions on a 5-point scale from (1) unrelated to (5) fully relevant and accurate (Amidei et al., 2019; Liu et al., 2023). Missing values were marked correct if the information was genuinely unavailable in the source papers. More details are in appendix B.
Table 3 shows strong extraction accuracy and description quality of our dataset. GPT-4o (Hurst et al., 2024), which we used during the data extraction, demonstrated effective long-context information extraction, showcasing its potential for scientific literature synthesis. Moreover, high validation scores suggest missing values stem from unreported setups. Cohen’s Kappa scores of 0.68 (extraction) and 0.57 (description) indicate substantial inter-annotator agreement (McHugh, 2012).
During the study, we recorded the time experts spent annotating target attribute information (excluding descriptions) from additional samples based on a list of papers and tables indexing experimental results of target models. On average, it took 7 minutes and 50 seconds per table, totaling approximately about 350 hours for our dataset of 2,694 tables. This does not include the initial effort of surveying and identifying papers and tables. In contrast, our pipeline identifies and extracts data from arXiv sources from 2023 to 2024 in a single day for under \$500 using a batching API.
@@ 3.2 Categorization by Required Skills
We categorize experimental records from our dataset by required skills to enable comprehensive meta-analyses and offer researchers a searchable resource. Expanding on Tulu3’s core skills (Lambert et al., 2024), we defined 10 categories: Knowledge, Reasoning, Math, Coding, Multimodality, Instruction Following, Safety, Multilinguality, Tool Use (Agent Framework), and Other. Records were classified using an LLM API (Hurst et al., 2024) based on dataset names, subsets, and descriptions. Since datasets can fit multiple categories, we applied multi-label classification, allowing multiple tags per record. This categorized information is used throughout the meta-analyses in section 4.1 and section 4.2.
@@
4 Meta Analysis of Prompting Behavior
Our extracted dataset enables partially automated meta-analyses of LLM prompting behaviors through its structured attributes. We validate our semi-automated approach to meta-analysis by replicating Sprague et al. (2024)’s manual analysis, confirming CoT’s advantages over direct prompting in mathematical and symbolic reasoning tasks (section 4.1). We then show that our dataset enables a fine-grained analysis, showing that in-context examples boost performance in multimodal tasks but not in mathematical ones (section 4.2). We qualitatively analyze dataset characteristics, such as required expert knowledge, that may affect the performance of CoT and ICL. Finally, we analyze CoT and ICL interactions, finding that while demonstrations enhance CoT performance, they do not affect CoT’s relative improvement over standard direct prompting (section 4.4).
@@ 4.1 Which Tasks Benefit from CoT?
CoT (Wei et al., 2022), a prompting technique for eliciting reasoning, has attracted considerable attention, with extensive literature on the subject (Wang et al., 2023; Yao et al., 2024) Recently, Sprague et al. (2024) conducted a manual meta-analysis, concluding that CoT improves over standard direct prompting primarily in mathematical and symbolic reasoning tasks. We extract instances from our dataset that meet the criteria for replicating their investigation.
This identification is achieved through simple filtering using the dataset’s attributes. We focus on instances from the same source paper and table, identifying experiments that feature both CoT and standard direct prompts under the same conditions (model, dataset, subset, metric, and few-shot setup). Human input is used only to identify whether a prompt is a CoT prompt or a standard direct prompt based on the prompting method attribute; this is the only manual effort required for the meta-analysis aside from implementing the analysis itself. We exclude CoT variations like "xx of Thoughts" (Yao et al., 2024) and CoT-SC (Wang et al., 2022). We then calculate CoT’s performance improvement over standard direct prompting, namely delta, where a positive delta indicates CoT outperforming direct prompting, and a negative delta indicates the reverse.
@@
Our analysis, shown in Fig.5, reveals that CoT significantly enhances performance in mathematical tasks, with both median and mean improvements surpassing the overall median. However, we do not see clear improvements for reasoning tasks, likely due to their complexity and diversity, such as commonsense and logical reasoning. To investigate further, we apply more detailed classification categories (but less general) from Sprague et al. (2024) and redo our study using LLM classification. The truncated results, presented in Fig. 6, clearly show that symbolic reasoning and mathematical tasks exhibit notable improvement over other reasoning and non-reasoning tasks. This confirms the reliability and efficiency of our dataset for meta-analyses, as it provides consistent results with those obtained through manual curation in previous studies. Statistical tests confirm these patterns, though performance varies considerably across different categories (std: 10.1). Refer to appendix C for more details.

### v2

# Can LLMs Help Uncover Insights about LLMs? A Large-Scale, Evolving Literature Analysis of Frontier LLMs

@@ Abstract
The surge of LLM studies makes synthesizing their findings challenging. Analysis of experimental results from literature can uncover important trends across studies, but the time-consuming nature of manual data extraction limits its use. Our study presents a semi-automated approach for literature analysis that accelerates data extraction using LLMs. It automatically identifies relevant arXiv papers, extracts experimental results and related attributes, and organizes them into a structured dataset, LlmEvalDB. We then conduct an automated literature analysis of frontier LLMs, reducing the effort of paper surveying and data extraction by more than 93% compared to manual approaches. We validate LlmEvalDB by showing that it reproduces key findings from a recent manual analysis of Chain-of-Thought (CoT) reasoning and also uncovers new insights that go beyond it, showing, for example, that in-context examples benefit coding & multimodal tasks but offer limited gains in math reasoning tasks compared to zero-shot CoT. Our automatically updatable dataset enables continuous tracking of target models by extracting evaluation studies as new data becomes available. Through LlmEvalDB and empirical analysis, we provide insights into LLMs while facilitating ongoing literature analyses of their behavior.
Can LLMs Help Uncover Insights about LLMs?
A Large-Scale, Evolving Literature Analysis of Frontier LLMs
@@ 1 Introduction
Figure 1: The diagram of our semi-automated literature analysis process and a key finding derived from data automatically extracted from the arXiv database.
@@
The rapid advancement of Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; Anthropic, 2024) has led to a proliferation of empirical studies. Recent “surveys of surveys” (Lv et al., 2024; Ruan et al., 2024) highlight the overwhelming growth of LLM publications, surpassing what individual researchers can manually review. The growing number of evaluations—each using different models, datasets, and prompting configurations—makes it challenging to analyze and synthesize findings across studies, underscoring the urgent need for automated analysis.
@@
A manual literature analysis that examines data from multiple independent studies to identify overall trends, offers a solution (Reiter, 2018; Hupkes et al., 2023). However, conducting such an analysis is time-consuming (Yun et al., 2023, 2024). For instance, as shown in Fig. 1, understanding the effects of combining Chain-of-Thought (CoT) with in-context examples requires identifying relevant papers, extracting specific experimental results, and aggregating the data. Furthermore, analyses based on existing studies are likely to become quickly outdated in a fast-moving field, motivating the need for automated analyses that can be continuously updated as new results are shared on preprint servers such as arXiv.org.
@@
In this study, we conduct the most extensive quantitative literature analysis of frontier LLMs to date. Our dataset, LlmEvalDB, comprises 18,127 experimental records extracted from 1,737 papers, along with 359 referenced papers describing datasets used in these experiments. LlmEvalDB can also be dynamically updated with minimal human effort to incorporate results from new models or studies as they are published. This comprehensive dataset enables us to present new insights into LLM performance, a first in the field where previous studies on extracting leaderboards from machine learning papers have primarily focused on improving data extraction accuracy (Kardas et al., 2020; Yun et al., 2024).
@@
To achieve this, we experiment with an automated approach to data extraction that efficiently processes research literature. This approach uses an LLM to scan the arXiv database, identify relevant papers with experimental results on target models, and extract experimental results along with pertinent attributes. This reduces the manual effort needed for surveying and extraction by more than 93% compared to expert manual data extraction. The approach employs schema-driven extraction (Bai et al., 2023) and context augmentation from paper contents to capture essential attributes for a comprehensive understanding of evaluation results. Additionally, the process creates dataset descriptions that summarize key characteristics of the dataset associated with the evaluation record, enhancing the dataset’s potential utility in literature analysis.
@@
Our analysis shows that LlmEvalDB can efficiently replicate a previous manual analysis (Sprague et al., 2024), which found that CoT reasoning primarily benefits mathematical and symbolic reasoning tasks. We also identify three novel insights from the dataset on prompting configurations. First, in-context learning (ICL) enhances coding & multimodal tasks but offers limited benefits for mathematical tasks compared to a zero-shot setup. Second, our qualitative analysis reveals specific characteristics of datasets that tend to reduce the effectiveness of CoT prompting and ICL. Third, we find that CoT prompting with demonstrations—commonly implemented through ICL—typically yields better performance than CoT without demonstrations. However, the relative improvement of CoT over standard direct prompting remains stable regardless of the number of demonstrations used (i.e., it holds true in both zero-shot and few-shot settings). We release our dataset and source code to support the ongoing automated integration of new findings in the literature.  [^1] The code is available at https://github.com/JJumSSu/meta-analysis-frontier-LLMs, and the dataset can be accessed at https://huggingface.co/datasets/jungsoopark/LLMEvalDB.
@@ 2 Data Extraction
Automated data extraction is essential for scaling and improving literature analysis efficiency. This section outlines our pipeline for extracting experimental results and model attributes from arXiv sources, which are used to construct LlmEvalDB. We define target models and attributes (section 2.1) and introduce our LLM-powered extraction process, which includes three stages: preprocessing & filtering (section 2.2), extraction & augmentation (section 2.3), and description generation (section 2.4). Unless stated otherwise, we utilize GPT-4o (Hurst et al., 2024) throughout the pipeline.
@@ 2.1 Defining Target Models and Attributes
To enable thorough analysis across different studies, we extract various performance-related fields from proprietary models. The targeted attributes include: Dataset Name, Subset, Model Name, Prompting Method, Number of Demonstrations, Metric Name, and Performance. Note that the subset refers to any division of the dataset, such as a subtask, domain, split, or language pair. For instance, as shown in Fig.2, the arXiv paper with ID 2301.08721 (Cheng et al., 2023) presents experimental results for GPT-4 (Achiam et al., 2023) on the SVAMP dataset (Patel et al., 2021). The evaluation used Batch Prompting with CoT (Wei et al., 2022; Cheng et al., 2023) and 12 in-context samples (Brown et al., 2020), achieving an accuracy score of 95.0.
@@ 2.2 Preprocessing & Filtering
To extract the experimental results of target models and attributes, we used arXiv sources from January 2023 to December 2024, focusing on machine learning papers (cs.CV, cs.AI, cs.CL, cs.LG). arXiv LaTeX sources preserve structure better than parsed PDFs, enabling easier dataset extraction—especially from tables—using simple regex, while PDFs require complex tools. Hence, following prior work (Kardas et al., 2020; Bai et al., 2023; Wang et al., 2024), we adopt arXiv LaTeX sources for data extraction, as over 90% of arXiv papers are submitted in LaTeX format (Frankston et al., 2024).
We downloaded LaTeX source files and applied regex-based methods to extract structured data like tables. Before extraction, we filtered tables to reduce time and API costs by selecting those with target model results. We also concentrated on leaderboard tables presenting benchmark results following Kardas et al. (2020) and used Llama3.1-70B-Instruct (Dubey et al., 2024) for the filtering.
@@ 2.3 Extraction & Augmentation
2.4 Dataset Description Generation
In addition to extracting structured representations of experiments, we also generate summary descriptions of the relevant tasks and datasets (Fig. 3) that can aid in-depth literature analysis, such as classifying records by research subarea. We use a two-stage approach to create these descriptions. Initially, the LLM generates descriptions using its internal knowledge based on the dataset name and subset, which is cost-effective. When the LLM is uncertain, especially with lesser-known datasets or due to a knowledge cut-off, we prompt it to extract descriptions using the full content of the source papers as a reference. The source papers can be either the paper containing the table or the original dataset paper cited within it, which is retrieved using the BibTex references obtained from section 2.3.
@@ 3 Comprehensive Analysis of the Dataset
Table 1 presents sampled instances of LlmEvalDB, while the dataset’s statistics are summarized in Table 2. A total of 18,127 experimental records of four target models were extracted from scanning over 300k arXiv source papers. GPT-4 (Achiam et al., 2023) and GPT-4o (Achiam et al., 2023) dominate the results, while Claude 3 Opus (Anthropic, 2024) and Gemini 1.0 Pro (Team et al., 2023) have significantly fewer entries, highlighting a bias favoring certain proprietary models. Many prompting configurations have missing values, confirmed by human verification (section 3.1) as unreported by the original authors, likely indicating the default setup. The majority of dataset descriptions are generated referencing the source paper of the table, as these papers often describe datasets when presenting benchmarks or detailing experimental sections.
3.1 Quality and Efficiency Assessment
Our human evaluation assessed extraction accuracy and description quality across 40 records of LlmEvalDB (280 total attributes) from different papers. Two NLP Ph.D. students verified field extraction correctness (Bai et al., 2023) and rated descriptions on a 5-point scale from (1) unrelated to (5) fully relevant and accurate (Amidei et al., 2019; Liu et al., 2023a). Missing values were marked correct if the information was genuinely unavailable in the source papers. More details are in appendix B.
Table 3 shows strong extraction accuracy and description quality of LlmEvalDB. GPT-4o (Hurst et al., 2024), which we used during the data extraction, demonstrated effective long-context information extraction, showcasing its potential for scientific literature synthesis. Moreover, high validation scores suggest missing values stem from unreported setups. Cohen’s Kappa scores of 0.68 (extraction) and 0.57 (description) indicate substantial inter-annotator agreement (McHugh, 2012).
During the study, we recorded the time experts spent annotating target attribute information (excluding descriptions) from additional samples based on a list of papers and tables indexing experimental results of target models. On average, it took 7 minutes and 50 seconds per table, totaling approximately about 350 hours for LlmEvalDB of 2,694 tables. This does not include the initial effort of surveying and identifying papers and tables. In contrast, our pipeline identifies and extracts data from arXiv sources from 2023 to 2024 in a single day for under \$500 using a batching API.
@@ 3.2 Categorization by Required Skills
We categorize experimental records from LlmEvalDB by required skills to enable comprehensive literature analyses and offer researchers a searchable resource. Expanding on Tulu3’s core skills (Lambert et al., 2024), we defined 10 categories: Knowledge, Reasoning, Math, Coding, Multimodality, Instruction Following, Safety, Multilinguality, Tool Use (Agent Framework), and Other. Records were classified using an LLM API (Hurst et al., 2024) based on dataset names, subsets, and descriptions. Since datasets can fit multiple categories, we applied multi-label classification. This categorized information is used throughout the analyses in section 4.
@@
4 Prompting Behavior in Frontier LLMs: A Literature Review
LlmEvalDB enables partially automated literature analyses of LLM-prompting behaviors through its structured attributes. We validate our semi-automated approach to literature analysis by replicating Sprague et al. (2024)’s manual analysis, confirming CoT’s advantages over direct prompting in mathematical and symbolic reasoning tasks (section 4.1). We then show that LlmEvalDB enables a fine-grained analysis, showing that in-context examples boost performance in coding & multimodal tasks but not in mathematical ones (section 4.2). We qualitatively analyze dataset characteristics, such as required expert knowledge, that may negatively affect the performance of CoT and ICL (section 4.3). Moreover, we analyze CoT and ICL interactions, finding that while demonstrations enhance CoT performance, they do not affect CoT’s relative improvement over standard direct prompting (section 4.4). Finally, we analyze a subset of peer-reviewed papers published in major journals or conferences, revealing consistent trends that further support the robustness and generalizability of our findings (section 4.5).
@@ 4.1 Which Tasks Benefit from CoT?
CoT (Wei et al., 2022), a prompting technique for eliciting reasoning, has attracted considerable attention, with extensive literature on the subject (Wang et al., 2023; Yao et al., 2024) Recently, Sprague et al. (2024) conducted a manual analysis, concluding that CoT improves over standard direct prompting primarily in mathematical and symbolic reasoning tasks. We extract instances from LlmEvalDB that meet the criteria for replicating their investigation.
This identification is achieved through filtering using the dataset’s attributes. We focus on instances from the same source paper and table, identifying experiments that feature both CoT and standard direct prompts under the same conditions (model, dataset, subset, metric, and few-shot setup). Human input is used only to identify whether a prompt is a CoT prompt or a standard direct prompt based on the prompting method attribute; this is the only manual effort required for the analysis aside from implementing the analysis itself. We exclude CoT variations like "xx of Thoughts" (Yao et al., 2024) and CoT-SC (Wang et al., 2022). We then calculate CoT’s performance improvement over standard direct prompting, namely delta, where a positive delta indicates CoT outperforming direct prompting, and a negative delta indicates the reverse.
@@
Our analysis, shown in Fig.5, reveals that CoT significantly enhances performance in mathematical tasks, with both median and mean improvements surpassing the overall median. However, we do not see clear improvements for reasoning tasks, likely due to their diversity, such as commonsense and logical reasoning.
To investigate further, we apply more detailed classification categories (but less general) from Sprague et al. (2024) and redo our study using LLM classification. The truncated results, presented in Fig. 6, clearly show that symbolic reasoning and mathematical tasks exhibit notable improvement over other reasoning and non-reasoning tasks. This confirms the reliability and efficiency of LlmEvalDB for literature analyses, as it provides consistent results with those obtained through manual curation in previous studies. Statistical tests, as shown in Total Results from Table. 7, confirm these patterns.